Assets 2012 2013
Cash 45 50
A/R 362 301
Inventory 386 237
Total 793 588
Net Fixed Assets 1000 1200
Total Assets 1793 1788
Liabilities 2012 2013
Accounts Payable 341 327
Notes Payable 150 125
Total 491 452
Long-Term Debt 500 750
Common Stock 400 400
Retained Earnings 95 125
1486 1727
2013
Sales 1553
Cost of Goods Sold 1242
Depreciation 200
EBIT 111
Interest 80
Taxes 60
Net Income -29

A. What is the Accounts Payable Turnover Ratio (to 1 decimal)?

B. What is the Inventory Period (to 1 decimal)?

C.What is the Accounts Receivable Period (to 1 decimal)?

D. What is the Operating Cycle (to 1 decimal)?
Answer #1

Accounts Payable Turnover Ratio=Cost of Goods Sold/Average Accounts Payable

Average Accounts Payable=(341+247)/2=$294

Accounts Payable Turnover Ratio=$1242/294=4.2 times

What is the Inventory Period=Accounts Payable Turnover Ratio/365 Days=365/4.25=85.9 days

What is the Accounts Receivable Period (to 1 decimal)?

=365/Account receivable turnover

Account receivable turnover =sales/average receivable

=1553/((362+301)/2)=4.7 times

Accounts Receivable Period =365/4.7=77.7

D. What is the Operating Cycle (to 1 decimal)?

Operating Cycle = Inventory Holding Period + Receivables Collection Period

=85.9+77.7=163.6 days